OpenStudy (anonymous):

An artist wants to make alabaster pyramids using a block of alabaster with a volume of 576 cubic inches. She plans to make each pyramid with a square base area of 3 square inches and a height of 4 inches. At most, how many pyramids can the artist make from the block of alabaster?

OpenStudy (anonymous):

Volume of Pyramid = area of base * height * 1/3 Volume of Pyramid = 3 * 4 * 1/3 = 4 ins^3 Theoretically the artist can make 576/4 =144 pyramids out of the block of alabaster!

OpenStudy (anonymous):

http://www.algebra.com/cgi-bin/Surface-area.faq.question.795311.html, that's where I got the answer x